OpenStudy (anonymous):

sec^2theta-tan^2theta=? explanation please

OpenStudy (anonymous):

sec^2(x) - tan^2(x) = 1/cos^2(x) - sin^2(x)/cos^2(x) = 1-sin^2(x)/cos^2(x) 1-sin^2(x) = cos^2(x) => from { cos^2(x) + sin^2(x) = 1 } = cos^2(x)/cos^2(x) sec^2(x) - tan^2(x) = 1

OpenStudy (anonymous):

Thankyou :)

OpenStudy (anonymous):

best of Luck

OpenStudy (anonymous):

sec(x) = 1/cos(x) ; tan(x) = sin(x) / cos(x) sin^2(x) + cos^2(x) =1 => cos^2(x) = 1 - sin^2(x)

OpenStudy (anonymous):

^ what did you just write? the answer is still 1 correct?

OpenStudy (anonymous):

nevermind i got it